# AIArena: A Blockchain-Based Decentralized AI Training Platform

### v1

@@ 1. Introduction
The growth of artificial intelligence (AI) has led to groundbreaking applications in many fields. However, it also brings challenges, especially because a few large companies control most of its development and use. This concentration of power increases bias in AI systems (challen2019artificial, ) and reduces public involvement in important decisions. This lack of transparency can exacerbate issues such as misuse of AI technologies or unethical practices, as there is little external oversight in the AI centralized systems. Such centralized control also slows down innovation in AI (Etzioni2016Democratizing, ; montes2019distributed, ) and leads to the unfair use of user data, which mainly benefits these companies. These compounded challenges underline the urgent need for a shift to decentralized AI (DeAI) (wang2024sok, ), where control and access are distributed across a broader and more inclusive range of participants.
Blockchain’s decentralized nature can help address the above challenges by allowing multiple participants to collaborate on AI development without relying on a central authority (chen2018machine, ; dong2024defending, ). Smart contracts (zou2019smart, ) can automate processes such as distributing rewards or verifying contributions, fostering fairness and efficiency. Blockchain also enhances data integrity, as participants can verify the source and ownership of data, preventing unauthorized use and ensuring compliance with ethical standards (salah2019blockchain, ).
@@ 2.4. Delegate Staking
Formally, the the reward for a delegator who delegates to a training node can be calculated as: $f_{i}\cdot\left(\sigma\cdot\frac{t_{d}}{t_{n}+t_{d}}\right)$, whereas $f_{i}$ refers to the total reward distributed to the training node $i$ and delegator based on the quality of the training node’s submission, $t_{d}$ is the stake amount from this given delegator, $t_{n}$ is the stake amount from the training node $i$. Similarly, the reward for a delegator who delegates to a validator can be calculated as: $\left(\sum_{i}{f_{i}(\Delta_{ji},s_{j})}\right)\cdot\sigma\cdot\frac{s_{d}}{s_{v}+S_{d}}$ in which $s_{d}$ refers to the stake amount from a given delegator and $s_{v}$ is the stake amount of the validator the delegator delegated to.
@@ 2.5. Various Validation Phases
    Challenging Phase: Spanning $x-x_{0}-x_{1}$ days, this phase allows any validator to issue a challenge if they suspect a model has been stolen. Training nodes must provide proof of legitimate training, such as proof of learning (jia2021proof, ; lan2021proof, ) or on-chain address-based watermarking (zhang2020model, ). If they fail to do so, the successful challenger receives the rewards allocated to the malicious training node, incentivizing honest participation.
@@
3. Implementation and Evaluation
3.1. On-Chain Implementation
Figure 2. Training and validation submissions over time.
@@ 3.2. Off-Chain Implementation
Code Co-creation (Task 14). Accurate code generation and understanding are critical when leveraging LLMs for advanced automation, particularly in low-resource blockchain languages. Building on this need, we focus our efforts on the Move language, aiming to develop a model capable of generating accurate instructions and comments. Through our AIArena platform, a diverse community of contributors collectively curated what we believe, to the best of our knowledge, to be the largest Move instruction dataset to date [^3] https://huggingface.co/datasets/flock-io/move-code-comment. Drawing from 514 distinct Move projects, they enriched the source code with detailed instructions and comments.

### v3

@@ Abstract.
Blockchain; Machine Learning; Decentralized Artificial Intelligence
journalyear: 2025copyright: ccconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3715484isbn: 979-8-4007-1331-6/2025/04ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learningccs: Theory of computation Design and analysis of algorithms
@@ 1. Introduction
The growth of artificial intelligence (AI) has led to groundbreaking applications in many fields. However, it also brings challenges, especially because a few large companies control most of its development and use. This concentration of power increases bias in AI systems (Challen et al., 2019) and reduces public involvement in important decisions. This lack of transparency can exacerbate issues such as misuse of AI technologies or unethical practices, as there is little external oversight in the AI centralized systems. Such centralized control also slows down innovation in AI (Etzioni, 2016; Montes and Goertzel, 2019) and leads to the unfair use of user data, which mainly benefits these companies. These compounded challenges underline the urgent need for a shift to decentralized AI (DeAI) (Wang et al., 2024), where control and access are distributed across a broader and more inclusive range of participants.
Blockchain’s decentralized nature can help address the above challenges by allowing multiple participants to collaborate on AI development without relying on a central authority (Chen et al., [n. d.]; Dong et al., 2024). Smart contracts (Zou et al., 2019) can automate processes such as distributing rewards or verifying contributions, fostering fairness and efficiency. Blockchain also enhances data integrity, as participants can verify the source and ownership of data, preventing unauthorized use and ensuring compliance with ethical standards (Salah et al., 2019).
@@ 2.4. Delegate Staking
Formally, the the reward for a delegator who delegates to a training node can be calculated as: $f_{i}\cdot(1-\sigma)\cdot\frac{t_{d}}{t_{n}+t_{d}}$, whereas $f_{i}$ refers to the total reward distributed to the training node $i$ and delegator based on the quality of the training node’s submission, $t_{d}$ is the stake amount from this given delegator, $t_{n}$ is the stake amount from the training node $i$. Similarly, the reward for a delegator who delegates to a validator can be calculated as: $\left(\sum_{i}{f_{i}(\Delta_{ji},s_{j})}\right)\cdot(1-\sigma)\cdot\frac{s_{d}}{s_{v}+s_{d}}$in which $s_{d}$ refers to the stake amount from a given delegator and $s_{v}$ is the stake amount of the validator the delegator delegated to.
@@ 2.5. Various Validation Phases
    Challenging Phase: Spanning $x-x_{0}-x_{1}$ days, this phase allows any validator to issue a challenge if they suspect a model has been stolen. Training nodes must provide proof of legitimate training, such as proof of learning (Jia et al., [n. d.]; Lan et al., 2021) or on-chain address-based watermarking (Zhang et al., 2020). If they fail to do so, the successful challenger receives the rewards allocated to the malicious training node, incentivizing honest participation.
@@
3. Implementation and Evaluation
3.1. On-Chain Implementation
Figure 2. Training and validation submissions over time.
@@ 3.2. Off-Chain Implementation
Code Co-creation (Task 14). Accurate code generation and understanding are critical when leveraging LLMs for advanced automation, particularly in low-resource blockchain languages. Building on this need, we focus our efforts on the Move language, aiming to develop a model capable of generating accurate instructions and comments. Through our AIArena platform, a diverse community of contributors collectively curated what we believe, to the best of our knowledge, to be the largest Move instruction dataset to date (see https://huggingface.co/datasets/flock-io/move-code-comment). Drawing from 514 distinct Move projects, they enriched the source code with detailed instructions and comments.
Acknowledgements
This work was partially supported by a research grant from the Ethereum Foundation.